OpenStudy (anonymous):

Find f ' (x) for the function. f (x)= (3x^5 -x^4)/ 15x

OpenStudy (anonymous):

\[f(x)=(3x^5)/15x -( x^4)/15x ---> f(x)=( x^4)/5 - (x^3)/15 \] f(x)' = \[(4x^3)/5 -( 3x^2)/15\]

OpenStudy (anonymous):

thank you:)